OpenStudy (madmerc):

Rationalize the denominator of square root of negative 49 over (7 - 2i) - (4 + 9i)

OpenStudy (madmerc):

the choices are \[\frac{ −49+2i }{−40 }\] \[\frac{ 49+21i }{ 58 }\] \[\frac{ 77+21i}{ -2 }\] \[\frac{ −77+21i }{ 130 }\]

OpenStudy (madmerc):

The problem looks like \[\frac{ \sqrt{-49} }{ (7 - 2i) - (4 + 9i) }\]

OpenStudy (campbell_st):

ok... so in the numerator use i^2 = -1 and also simplify the denominator by collecting like terms \[\frac{\sqrt{49i^2}}{3 - 11i}\] now multiply by the conjugate pair \[\frac{7i}{3 - 11i} \times \frac{3 + 11i}{3 + 11i}\] do the multiplication and then simplify

OpenStudy (madmerc):

How do I multiply them? Like straight across or

OpenStudy (anonymous):

The answer is \[\frac{ -77+21i }{ 130}\]

OpenStudy (anonymous):

I showed the way to solve in the other question

OpenStudy (campbell_st):

why give an answer.... how does that help understanding... ? and yes you multiply straight across, just like multiplying any 2 fractions \[\frac{7i(3 + 11i)}{(3 - 11i)(3 + 11i)}\] you should recognise the denominator is the difference of 2 squares... hope it helps